OpenStudy (anonymous):

If the average of 12 consecutive odd integers is 328, what is the least of these integers?

OpenStudy (saifoo.khan):

k + (k+2) + (k+4)... = 328

OpenStudy (anonymous):

Figured it out! Yayy, but sort of really manually. Here's what I do: 12 consecutive odd numbers right? and 328 is the average. so 325 and 327 are going to be the mid values. I count back 5 odd numbers from 325 that is 317! and voila that's teh correct answer! If anyone knows any other quick solution or formula thingy let me know please!

OpenStudy (apoorvk):

That's the best way actually, and if I were in your place, would have done the same thing (I don't remember formulae that well) But if you do need a formula, here's what we can do: Let the first (and the least) odd no. be 2n+1 ...(since 2n+1is always odd) So, the next 11 will be --> (2n+3, 2n+5, 2n+7,....2n+23) all the 12 sum up 24n+144 now, 24n+144 ------- = 328 12 or, 24n+144 = 12*328 = 3936 or, n = 3792/24 =158 so, the first odd no. = 2n+1 = 317 Same answer, albeit through a longer, time-consuming and rote process. But comes in handy just in case your brain isn't working.